OpenStudy (anonymous):

A bus decreases its speed from 80 km/h in 5s find acceleration of the bus

OpenStudy (anonymous):

Initial velocity(u)=80 km/h =5/18 m/s = 1km/h 80km/h=80 x 5/18=22.22m/s Final velolity (v)=60km/h =5/18m/s=1km/h 60km/h=60 x 5/18=16.66m/s t=5s a=v-u/t a=16.66-22.22/5 =12.216 m/s2

OpenStudy (shane_b):

The question isn't clear to me. The bus decreased its speed from 80 km/h to what? 60 km/h as you wrote? Or is that supposed to mean that it decreased its speed BY 80km/h? If it went from 80km/h to 60km/h then the difference in velocity is 20km/h: \[\frac{20km}{1hr}=\frac{20000m}{3600s}=5.56m/s\]\[a=\frac{\Delta V}{t}=\frac{-5.56m/s}{5s}=-1.11m/s^2\] If it decreased its speed BY 80 km/h the difference in velocity is obviously 80km/h: \[\frac{80km}{1hr}=\frac{80000m}{3600s}=22.22m/s\]\[a=\frac{\Delta V}{t}=\frac{-22.22m/s}{5s}=-4.444m/s^2\]

OpenStudy (shane_b):

Don't forget to adjust for significant digits as well